OpenStudy (anonymous):

Does the point (2,3) lie on the graph y=3x-4? Explain

OpenStudy (kat_s123):

Plug the points in for x and y. So 3=3(2)-4 and then simplify. 3=6-4. 3 does not equal 2. So the point does not fall on the line.